OpenStudy (anonymous):

A public library has fiction and nonfiction books in the ratio three colon two. If the library has three hundred sixty fiction books, how many nonfiction books does it have? seventy-two one hundred twenty one hundred eighty two hundred forty

OpenStudy (anonymous):

can u show ur work pls

OpenStudy (anonymous):

@texaschic101 @Wolfboy @WordGEEK @iambatman @paki

OpenStudy (texaschic101):

fiction 3 360 ------- = ---- = ---- non-fiction 2 x this is a proportion, so cross multiply and solve for x. (3)(x) = (360)(2) 3x = 720 x = 720/3 x = 240 check.. 3/2 = 360/240 cross multiply (2)(360) = (3)(240) 720 = 720 (correct) if there are 360 fiction, then there are 240 non-fiction

OpenStudy (anonymous):

thanks!!